OpenStudy (anonymous):

A certain forest covers an area of 5000km^2 . Suppose that each year this area decreases by 5.75% . What will the area be after 14 years? the formula eludes me...I don't trust mine calcultion Use the calculator provided and round your answer to the nearest square kilometer.

OpenStudy (anonymous):

5000-(5000*.0575)=x n=1 x-x(.0575)=x2 n=2 ... x(n-1)-(x(n-1)*.0575)=xn Continue until x14